OpenStudy (anonymous):

Kenya exchanges $200 for euros (€). Suppose the conversion rate is €1 = $1.321. How many euros should Kenya receive?

OpenStudy (anonymous):

set up a proportion.. 1/1.321 = 200/x (1 euro to $1.321 = 200 euros to x dollars) cross multiply (1)(x) = (1.321)(200) can you finish and solve for x , the numbers of dollars ?

OpenStudy (anonymous):

is it 264.20

OpenStudy (anonymous):

yes...you got it . $264.20

OpenStudy (anonymous):

good job :)

OpenStudy (anonymous):

thank you!